OpenStudy (anonymous):

If ADB = 70°, then ABD = degrees m AB = degrees m AD = degrees

OpenStudy (amistre64):

if all you know is one degrees; ... this has got to have a picture attached

OpenStudy (amistre64):

:) i learned to day after much consternation, that the tangent is half the arc :)

OpenStudy (saifoo.khan):

ABD = 20degrees

OpenStudy (saifoo.khan):

90 + 70 + x =180 x = 20*

OpenStudy (amistre64):

i dont think that pic is to scale ....

OpenStudy (anonymous):

AB = 140 degrees AD = 40 degrees

OpenStudy (anonymous):

This is not hard at all.

OpenStudy (anonymous):

Since ADB is an inscribed angle, the arc it cuts will be twice the measure of the angle. Arc BD = 180 degrees. Arc BD = Arc BA + Arc AD = 140 + 40 = 180 degrees.

OpenStudy (anonymous):

terricherri, this doesn't work well without your feedback...